OpenStudy (anonymous):

If f(x) = 2x2 − 30, find f(4).

OpenStudy (eyust707):

Change all the "x"s to "4"s

hartnn (hartnn):

put x=4 what u got ?

OpenStudy (anonymous):

that is 2^2

OpenStudy (anonymous):

it would be 34?

hartnn (hartnn):

nopes, 2(4^2)-30 = 2* 16-30 = ??

OpenStudy (anonymous):

2

OpenStudy (anonymous):

is that right?

OpenStudy (eyust707):

Yes

OpenStudy (anonymous):

thanks